# Data-based control of Logical Networks

### v1

@@ 5 Data-driven output regulation
    $\left[\begin{array}[]{c|c}I_{N}&\mathbbold{0}_{N\times P}\cr\hline\cr\mathbbold{0}_{1\times N}&({\bf y}^{*})^{\top}\end{array}\right]\left[\begin{array}[]{c}X_{p}\cr\hline\cr Y_{p}\end{array}\right]=\left[\begin{array}[]{c}X_{p}\cr\hline\cr{\bf v}^{\top}\end{array}\right]$
    Note that ${\bf v}^{\top}=({\bf y}^{*})^{\top}Y_{p}\in\mathcal{B}^{T}$, and $[{\bf v}]_{k}\neq 0$ if and only if ${\rm col}_{k}(X_{p})\in\mathcal{X}^{d}({\bf y}^{*})$.
  • Define $\mathcal{K}\triangleq\{k_{1},k_{2},\dots,k_{y}\}=$nonzero(${\bf v}$)
@@
            if $\delta_{N}^{k}\notin\mathcal{X}^{d}({\bf y}^{*})$, then
                $\mathcal{X}^{d}({\bf y}^{*})\leftarrow\mathcal{X}^{d}({\bf y}^{*})\cup\{\delta_{N}^{k}\}$
@@
Conclusion: Return $\mathcal{X}^{d}({\bf y}^{*})$. ∎
At this point, we need to verify if $\mathcal{X}^{d}({\bf y}^{*})$ contains some limit cycles. It is worth highlighting that to determine limit cycles of unitary length we can use Algorithm 1. However, the next algorithm is able to also identify equilibrium points, being them a special case of limit cycles. To this end, let us define the data matrix
$X\!\triangleq\!\begin{bmatrix}{\bf x}_{d}(0)&{\bf x}_{d}(1)&\dots&{\bf x}_{d}(T-1)&{\bf x}_{d}(T)\end{bmatrix}\in\mathcal{L}_{N\times(T+1)},$
containing all the logical vectors of the sequence ${\bf x}_{d}$. The first step consists in determining all the transitions appearing in the matrix $X$ involving (both initial and final) states that belong to $\mathcal{X}^{d}({\bf y}^{*})$. Based on these transitions, we then construct a digraph, whose set of nodes is $\mathcal{X}^{d}({\bf y}^{*})$, and with edges corresponding to the transitions identified from the data. We call $\tilde{L}_{tot}$ the Boolean (but not necessarily logical) matrix associated to such a digraph, since we are accounting for transitions associated to different inputs. The matrix $\tilde{L}_{tot}$ represents the data-based estimate of the principal submatrix of $L_{tot}$ in (7) obtained by selecting the rows and the columns corresponding to the states in $\mathcal{X}^{d}({\bf y}^{*})$. Indeed, as the data are not informative for identifiability of the BCN, the unitary entries of $\tilde{L}_{tot}$ in general are only a subset of those of $L_{tot}$ for the original BCN (and for every BCN in ${\mathcal{B}}_{d}^{0}$). We now exploit the well-know Johnson’s Algorithm Johnson (1975) to identify all limit cycles in ${\mathcal{D}}(\tilde{L}_{tot})$.
Algorithm 5333The proposed algorithm is tailored for a sequence of data collected in a single experiment. If the performed experiments are more than one, it must be applied separately to each experiment, and then the results must be integrated. : Limit cycles in $\mathcal{X}^{d}({\bf y}^{*})$
@@
           - The set $\mathcal{X}^{d}({\bf y}^{*})$.
Output: The set $\mathcal{C}^{d}({\bf y}^{*})$ of all limit cycles in $\mathcal{X}^{d}({\bf y}^{*})$.
Initialization:
Set:
  • $P\triangleq\begin{bmatrix}\delta_{N}^{i_{1}}&\delta_{N}^{i_{2}}&\dots&\delta_{N}^{i_{s}}\end{bmatrix}\in{\mathcal{L}}_{N\times s},$ where $s$ is the cardinality of $\mathcal{X}^{d}({\bf y}^{*})$ and $\mathcal{I}\triangleq\{i_{1},i_{2},\dots,i_{s}\}\subseteq[1,N]$ is the set of indices of the states in $\mathcal{X}^{d}({\bf y}^{*})$ (taken in an arbitrary order);
@@
  • $\mathcal{D}(\tilde{L}_{tot})\triangleq(\tilde{\mathcal{V}},\tilde{\mathcal{E}})$, with $\tilde{\mathcal{V}}=\mathcal{X}^{d}({\bf y}^{*})$ and $\tilde{\mathcal{E}}=\emptyset$.
Procedure:
if $\mathcal{I}\neq\emptyset$, then
  • Compute $X^{*}\triangleq P^{\top}X\in\mathcal{L}_{s\times(T+1)}$
  • for $(i,j)\in[1,s]\times[1,T+1]$, do
          if $\exists k\in[1,T]$ s.t. $[X^{*}]_{ik}[X^{*}]_{j,k+1}\neq 0$, then
              ~~$\tilde{\mathcal{E}}\leftarrow\tilde{\mathcal{E}}\cup\{(i,j)\}$
@@
  • Apply Johnson’s Algorithm in Johnson (1975) to find all cycles in $\mathcal{D}(\tilde{L}_{tot})$, say $\tilde{\mathcal{C}}_{1},\dots,\tilde{\mathcal{C}}_{n_{c}}$, $n_{c}\in\mathbb{Z}_{+}$
  • Remap all the nodes of $\mathcal{D}(\tilde{L}_{tot})$ to the corresponding nodes in $[1,N]$, by using the bijective correspondence given by the matrix $P$, i.e., $1\leftrightarrow i_{1},\dots,s\leftrightarrow i_{s}$
  • Define $\mathcal{C}_{1},\dots,\mathcal{C}_{n_{c}}$ the remapped limit cycles
  • $\mathcal{C}^{d}({\bf y}^{*})=\{\mathcal{C}_{1},\dots,\mathcal{C}_{n_{c}}\}$
else Go to the Conclusion
end if
Conclusion: Return $\mathcal{C}^{d}({\bf y}^{*})$. ∎
For each edge $(i,j)$ belonging to some limit cycle ${\mathcal{C}}_{l}$ in $\mathcal{C}^{d}({\bf y}^{*})$ it is also possible to determine one of the inputs related to such transition within the limit cycle, say ${\bf u}_{i}^{l}$. Indeed, if there exists an edge $(i,j)$, it means that the transition $i\to j$ must be captured by the collected data, or, equivalently, there exists $k\in[1,T]$ such that $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{i}\cr\delta_{N}^{j}\end{bmatrix}.$ Therefore, the input associated to this transition within ${\mathcal{C}}_{l}$ is ${\bf u}_{i}^{l}\triangleq U_{p}\delta_{T}^{k}.$ At this point it only remains to establish if the set $\mathcal{C}^{d}({\bf y}^{*})$ is globally reachable. This is possible by applying (a slightly amended version of) Algorithm 3.
We can now provide the solution to Problem 3.
@@
Problem 3 is solvable (i.e., the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for output regulation to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback) if and only if the following two conditions hold:
  • (i)
@@
  • (ii)
    ${\mathcal{S}^{*}_{o}}\triangleq\bigcup_{{\bf x}\in\mathcal{C}^{d}({\bf y}^{*})}\mathcal{S}({\bf x})=\mathcal{L}_{N}$.
Moreover, the generic $r$-th column of a possible feedback matrix $K\in{\mathcal{L}}_{M\times N}$ is given by ${\rm col}_{r}(K)={\bf u}_{r},\forall r\in[1,N],$ where ${\bf u}_{r}$ is determined by Algorithm 3.
@@
(Necessity). Assume, by contradiction, that condition (i) does not hold, i.e., $\mathcal{C}^{d}({\bf y}^{*})=\emptyset$. But this means that there exists at least one BCN in $\mathcal{B}_{d}$ for which $\mathcal{C}({\bf y}^{*})=\emptyset$, and hence, by Proposition 18, for that BCN the output regulation to ${\bf y}^{*}$ by state feedback is not possible.
Assume now that $\mathcal{C}^{d}({\bf y}^{*})\neq\emptyset$, but (ii) does not hold, namely $\mathcal{S}^{*}\neq\mathcal{L}_{N}$. We can construct a BCN in ${\mathcal{B}}_{d}^{o}$ for which $\mathcal{C}({\bf y}^{*})=\mathcal{C}^{d}({\bf y}^{*})$ and $\bigcup_{{\bf x}\in\mathcal{C}({\bf y}^{*})}\mathcal{S}({\bf x})={\mathcal{S}^{*}_{o}}\neq\mathcal{L}_{N}$. This requires to impose all the state transitions, equivalently the columns of $\tilde{L}$, that are not uniquely identified by the data (and possibly also the corresponding output). If there exists $\delta^{r}_{N}$ that does not appear as a column of $X_{p}$, then for every $i\in[1,M]$ we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ and $H\delta^{r}_{N}\neq{\bf y}^{*}$. On the other hand, if $\delta^{r}_{N}=X_{p}\delta^{k}_{T}$ for some $k\in[1,T]$, we define $\mathcal{I}_{r}$ the set of indices $i\in[1,M]$ such that $U_{p}\delta^{k}_{T}=\delta^{i}_{M}$. If $\mathcal{I}_{r}\subsetneq[1,M]$, we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ for every $i\not\in\mathcal{I}_{r}$. Clearly, by Proposition 18, for each BCN obtained in this way the output regulation to ${\bf y}^{*}$ by state feedback is not possible.
(Sufficiency). It is a direct consequence of Proposition 18 and following comments.
The final part follows trivially from Algorithm 3. ∎
@@
Assume that we want to regulate the output of the network to the value ${\bf y}^{*}=\delta_{2}^{2}$. We perform an offline experiment in the time interval $[0,9]$, and collect the data:
@@
where, again, the symbol * stands for an arbitrary vector (in ${\mathcal{L}}_{6}$). As $X_{p}$ is of full row rank, the matrix $H$ can be uniquely recovered. Hence, by applying Algorithm 4, we obtain that the set of all the limit cycles in $\mathcal{X}^{d}({\bf y}^{*})$ is $\mathcal{C}^{d}({\bf y}^{*})=\{\delta_{6}^{3},\delta_{6}^{2}\leftrightarrow\delta_{6}^{4}\},$ where $\delta_{6}^{2}\leftrightarrow\delta_{6}^{4}$ is the limit cycle of length $2$ involving $\delta_{6}^{2}$ and $\delta_{6}^{4}$. We now check if the set $\mathcal{C}^{d}({\bf y}^{*})$ is globally reachable. By following the procedure outlined in Algorithm 3, we obtain $\mathcal{S}_{o}^{*}={\mathcal{L}}_{6}$. Thus, by Theorem 19, the data are informative for output regulation to ${\bf y}^{*}$ by state feedback. The feedback matrix $K=\begin{bmatrix}\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}\end{bmatrix}$ is obtained via Algorithm 3.
@@ 6 Conclusion
In this paper we have investigated how the recently introduced data informativity approach van Waarde et al. (2020) can be adapted and extended to solve fundamental control problems in Boolean control networks. More specifically, we focused on the feedback stabilization, safe control and output regulation problems, and we provided necessary and sufficient conditions, based only on previously collected data (supposed not to be informative for identification), for their solvability. Moreover, we also obtained from data the feedback matrices expressions. We validated our results by means of examples.

### v2

@@ 7 Data-driven output regulation
    ${\bf v}^{\top}=({\bf y}^{*})^{\top}Y_{p}\in\mathcal{B}^{1\times T},$
    and note that ${\bf v}]_{k}\neq 0$ if and only if $X_{p}\delta_{T}^{k}\in\mathcal{X}^{d}({\bf y}^{*})$.
  2. 2.
    Define $\mathcal{K}\triangleq\{k_{1},k_{2},\dots,k_{y}\}=\$nonzero($\bf v$})
  3. 3.
@@
        if $X_{p}\delta_{T}^{k}\notin\mathcal{X}^{d}({\bf y}^{*})$, then
          $\mathcal{X}^{d}({\bf y}^{*})\leftarrow\mathcal{X}^{d}({\bf y}^{*})\cup\{X_{p}\delta_{T}^{k}\}$
@@
  4. 4.
    Return $\mathcal{X}^{d}({\bf y}^{*})$.
At this point, we need to verify if there exist cycles whose nodes belong to $\mathcal{X}^{d}({\bf y}^{*})$. The first step consists in determining all the transitions appearing in the data involving (both initial and final) states that belong to $\mathcal{X}^{d}({\bf y}^{*})$. Based on these transitions, we then construct a digraph, whose set of nodes is $\mathcal{X}^{d}({\bf y}^{*})$, and with edges corresponding to the transitions within $\mathcal{X}^{d}({\bf y}^{*})$ identified from data. We call $L_{tot}^{d}({\bf y}^{*})$ the Boolean (but not necessarily logical) matrix associated to such a digraph. The matrix $L_{tot}^{d}({\bf y}^{*})$ represents the data-based estimate of the principal submatrix of $L_{tot}$ in (15) obtained by selecting the rows and the columns corresponding to the states in $\mathcal{X}^{d}({\bf y}^{*})$. In Algorithm 4, we exploit the well-know Johnson’s Algorithm (see Johnson (1975)) to identify the set of all cycles in ${\mathcal{D}}(L_{tot}^{d}({\bf y}^{*}))$, say $\mathcal{C}^{d}({\bf y}^{*})$.
Algorithm 4 Cycles in $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$
@@
     - The set $\mathcal{X}^{d}({\bf y}^{*})=\{\delta^{i_{1}}_{N},\dots,\delta^{i_{s}}_{N}\}$.
Output: The set $\mathcal{C}^{d}({\bf y}^{*})$ of all cycles in $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$.
Initialization: Set:
  • $P\triangleq\begin{bmatrix}\delta_{N}^{i_{1}}&\delta_{N}^{i_{2}}&\dots&\delta_{N}^{i_{s}}\end{bmatrix}\in{\mathcal{L}}_{N\times s}$;
  • $\tilde{\mathcal{V}}\triangleq\{i_{1},i_{2},\dots,i_{s}\}\subset[1,N]$;
@@
[…]
  1. 1.
    for $(i,j)\in[1,s]\times[1,s]$, do
    if $\exists k\in[1,T]$ s.t. $P^{\top}X_{p}]_{ik}[P^{\top}X_{f}]_{jk}\neq 0$, then
         $\tilde{\mathcal{E}}\leftarrow\tilde{\mathcal{E}}\cup\{(i,j)\}$
@@
  2. 2.
    Apply Johnson’s Algorithm in Johnson (1975) to find all cycles in $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$, say $\tilde{\mathcal{C}}_{1},\dots,\tilde{\mathcal{C}}_{n_{c}}$, $n_{c}\in\mathbb{Z}_{+}$.
  3. 3.
    Remap all the nodes of $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$ to the corresponding nodes in $[1,N]$, by using the bijective correspondence given by the matrix $P$, i.e., $1\leftrightarrow i_{1},\dots,s\leftrightarrow i_{s}$.
  4. 4.
    Define $\mathcal{C}_{1},\dots,\mathcal{C}_{n_{c}}$ the remapped limit cycles.
  5. 5.
    Return $\mathcal{C}^{d}({\bf y}^{*})$.
For each edge $(i,j)$ belonging to some cycle $\mathcal{C}$$_{\ell}$ in $\mathcal{C}^{d}({\bf y}^{*})$ it is also possible to determine one of the inputs associated with such transition within the cycle, say ${\bf u}_{i}^{\ell}$. Indeed, if there exists an edge $(i,j)$, it means that the transition $i\to j$ is captured by the collected data, or, equivalently, there exists $k\in[1,T]$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{i}\cr\delta_{N}^{j}\end{bmatrix}$. (26)
Therefore, the input associated to this transition within $\mathcal{C}$$_{\ell}$ is
${\bf u}_{i}^{\ell}\triangleq U_{p}\delta_{T}^{k}.$ (27)
At this point it only remains to establish if the set $\mathcal{C}^{d}({\bf y}^{*})$ is globally reachable. This can be verified using Algorithm 1, which additionally specifies the inputs to be applied to the states outside $\mathcal{C}^{d}({\bf y}^{*})$.
We can now provide the solution to Problem 2.
@@
Problem 2 is solvable (i.e., the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for output regulation to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback) if and only if the following two conditions hold:
  • i)
@@
  • ii)
    ${\mathcal{S}^{*}}\triangleq\bigcup_{{\bf x}\in\mathcal{C}^{d}({\bf y}^{*})}\mathcal{S}({\bf x})=\mathcal{L}_{N}$.
@@
(Necessity). Assume, by contradiction, that condition i) does not hold, i.e., $\mathcal{C}^{d}({\bf y}^{*})=\emptyset$. But this means that there exists at least one BCN in $\mathcal{B}_{d}$ for which $\mathcal{C}({\bf y}^{*})=\emptyset$, and hence, by Proposition 15, for that BCN the output regulation to $\bf y$$^{*}$ by state feedback is not possible.
Assume now that $\mathcal{C}^{d}({\bf y}^{*})\neq\emptyset$, but ii) does not hold, namely $\mathcal{S}^{*}\neq\mathcal{L}_{N}$. We can construct a BCN in ${\mathcal{B}}_{d}$ for which $\mathcal{C}({\bf y}^{*})=\mathcal{C}^{d}({\bf y}^{*})$ and $\bigcup_{{\bf x}\in\mathcal{C}({\bf y}^{*})}\mathcal{S}({\bf x})={\mathcal{S}^{*}}\neq\mathcal{L}_{N}$. This requires to select all the state transitions, equivalently the columns of $\tilde{L}$, that are not uniquely identified by the data (and possibly also the corresponding output). If there exists $\delta^{r}_{N}$ that does not appear as a column of $X_{p}$, then for every $i\in[1,M]$ we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ and $H\delta^{r}_{N}\neq{\bf y}^{*}$. On the other hand, if $\delta^{r}_{N}=X_{p}\delta^{k}_{T}$ for some $k\in[1,T]$, we define $\mathcal{I}_{r}$ the set of indices $i\in[1,M]$ such that $U_{p}\delta^{k}_{T}=\delta^{i}_{M}$. If $\mathcal{I}_{r}\subsetneq[1,M]$, we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ for every $i\not\in\mathcal{I}_{r}$. Clearly, by Proposition 15, for each BCN obtained in this way the output regulation to $\bf y$$^{*}$ by state feedback is not possible.
(Sufficiency). It is a direct consequence of Proposition 15 and following comments. ∎
As for Theorem 12 in Section 6, Theorem 16 only provides a method to verify, using data, whether the output regulation problem is solvable. However, it is not constructive, in the sense that it does not offer a systematic procedure for explicitly computing the solution, namely the feedback matrix $K\in\mathcal{L}_{M\times N}$. Nevertheless, the matrix $K$ can be readily obtained as
$K=\begin{bmatrix}{\bf u}_{1}&\dots&{\bf u}_{N}\end{bmatrix}$
by selecting as control input ${\bf u}_{i}$ the one defined in (27) for the states $\delta^{i}_{N}$ belonging to a cycle in $\mathcal{C}^{d}({\bf y}^{*})$, and using the input ${\bf u}_{i}$ provided by Algorithm 1 for all other states $\delta^{i}_{N}$. In cases where a state belongs to multiple cycles in $\mathcal{C}^{d}({\bf y}^{*})$, it suffices to select the control input associated with any of these cycles.
Remark 17.
In this section, we have explored how to solve the output regulation problem via state feedback based on collected data. Clearly, this strategy cannot be implemented unless offline data about the state have been collected. An alternative approach, first explored in Leifeld et al. (2018), is to perform output regulation by means of output feedback. When so, one could try to achieve this goal by collecting only input and output data. The solution proposed in Leifeld et al. (2018), and inspired by results on the stabilization of Probabilistic Boolean networks, relies on an algorithm that first computes an approximation of the output prediction matrix, based on data, and then generates a sequence of concentric annular sets, each containing the output values that lie at a specific “distance" from the desired value $\bf y$$^{*}$, with respect to the output/input/output transitions captured by the data-based output prediction matrix. Finally, using these sets, the algorithm computes, when possible, an output feedback gain.
However, output feedback strategies, which are demanding even in model-based settings, become hardly applicable in a data-driven framework, where only part of the network transitions is accessible.
@@
Assume that we want to regulate the output of the network to the value ${\bf y}^{*}=\delta_{2}^{2}$. We perform a single ($r=1$) offline experiment in the time interval $[0,9]$, and collect the data:
@@
where, again, the symbol * stands for an arbitrary vector (in ${\mathcal{L}}_{6}$). As $X_{p}$ is of full row rank, the matrix $H$ can be uniquely recovered. Hence, by applying Algorithm 4, we obtain that the set of all cycles in $\mathcal{D}(L_{tot}^{d}({\bf y}^{*}))$ is $\mathcal{C}^{d}({\bf y}^{*})=\{3,2\leftrightarrow 4\},$ where $2\leftrightarrow 4$ is the cycle of length $2$ involving $\delta_{6}^{2}$ and $\delta_{6}^{4}$. We now check if the set $\mathcal{C}^{d}({\bf y}^{*})$ is globally reachable. By following the procedure outlined in Algorithm 1, we obtain $\mathcal{S}^{*}={\mathcal{L}}_{6}$. Thus, by Theorem 16, the data are informative for output regulation to $\bf y$$^{*}$ by state feedback. Building on the argument presented after Theorem 16, we obtain the following feedback matrix
$K=\begin{bmatrix}\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{2}&\delta_{3}^{3}&\delta_{3}^{2}\end{bmatrix}$.
@@ 8 Conclusion
In this paper we investigated how the recently introduced data informativity approach van Waarde et al. (2020) can be adapted and extended to solve fundamental control problems in Boolean control networks. More specifically, we focused on safe control and output regulation problems, and we provided necessary and sufficient conditions, based only on previously collected data (supposed not to be informative for identification), for their solvability. Moreover, we also obtained from data the feedback matrices expressions. We validated our results by means of examples. It is worth noting that, due to the inherent complexity of these systems, BCN problems are NP-hard with respect to the state, input and output dimensions $n$, $m$ and $p$, or equivalently polynomial in the corresponding quantities $N=2^{n}$, $M=2^{m}$ and $P=2^{p}$. Importantly, the computational complexity is independent of the chosen representation: in the case of interest, the algebraic one. Moreover, relying on data instead of an exact model does not substantially change the overall computational complexity. The only difference is that it also depends on the collected data, and in particular on the number of collected data $T$, which is, however, lower-bounded by $N$. On the other hand, since data provide only partial knowledge of the system dynamics, the number of identified transitions is smaller than those appearing in the real model, and consequently, if their complete enumeration is required, it imposes a lower computational burden. Nonetheless, depending on the problem at hand, the algorithms can be easily adapted to terminate once a single solution is found, without the need to search for all feasible solutions.